# Multi-Class Stackelberg Games for the Co-Design of Networked Systems

### v1

@@ III-B2 Class II: Stackelberg game with cooperative leaders and cooperative followers
Similar to the previous class, both sets of decision-makers, i.e., leaders $\mathcal{L}$ and $\mathcal{M}$, behave in the same strategic manner. In this combination of strategic actions, both sets across the layers cooperate within their corresponding layer. This structure leads to the computation of a solution for centralized optimization problems. The following are the steps to compute the Stackelberg equilibrium under this game class:
  • For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the optimal solution for the cooperative problem given by $\tilde{\mathbf{u}}^{*}\in\mathcal{U}$, which is denoted by $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ as it depends on the strategic selection of the leader. The computed optimal is a cooperative best response against the leaders’ strategies, i.e., $\tilde{\mathbf{u}}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{C}}_{\mathcal{M}}(\mathbf{a})$.
  • Using the optimal strategic profile $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader cooperative game problem can be stated only in terms of $\mathbf{a}$, and the optimal solution for the leaders $\tilde{\mathbf{a}}^{*}\in\mathcal{A}$ can be found.
  • Then, using the optimal solution for the leaders game problem $\tilde{\mathbf{a}}^{*}$, the cooperative optimal solution for the follower becomes $\tilde{\mathbf{u}}^{*}:=\tilde{\mathbf{u}}^{*}(\tilde{\mathbf{a}}^{*})$.
  • Finally, one obtains the Class II Stackelberg equilibrium as $(\tilde{\mathbf{u}}^{*},\tilde{\mathbf{a}}^{*})\in\mathcal{U}\times\mathcal{A}$.
Algorithm III-B2 shows the detailed procedure to compute the Stackelberg equilibrium for this class.
@@ III-B3 Class III: Stackelberg game with non-cooperative leaders and cooperative followers
This class comprises a heterogeneous strategic behavior for the leaders $\mathcal{L}$ with respect to the followers $\mathcal{M}$. In this class, the leaders behave selfishly as they independently optimize a cost function, i.e., this layer is solved by following a best-response dynamics to compute the corresponding solution in the normal form game. Then, as the followers jointly optimize to compute the control input, this step consists of a standard centralized optimal control problem. Next, we present the steps to compute the Stackelberg equilibrium for this class.
  • For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the optimal solution for the cooperative problem that corresponds to a control problem given by $\tilde{\mathbf{u}}^{*}\in\mathcal{U}$, which can be denoted by $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ as it depends on the strategic selection of the leader. The computed optimal is a cooperative best response against the leaders’ strategies, i.e., $\tilde{\mathbf{u}}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{C}}_{\mathcal{M}}(\mathbf{a})$.
  • Using the optimal strategic profile $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader cooperative game problem can be stated only in terms of $\mathbf{a}$, and the Nash equilibrium for the leaders $\mathbf{a}^{*}\in\mathcal{A}$ can be found.
  • Then, using the Nash equilibrium solution for the leaders game problem $\mathbf{a}^{*}$, the cooperative optimal solution for the follower becomes $\tilde{\mathbf{u}}^{*}:=\tilde{\mathbf{u}}^{*}(\mathbf{a}^{*})$.
  • Finally, one obtains the Class III Stackelberg equilibrium as $(\tilde{\mathbf{u}}^{*},\mathbf{a}^{*})\in\mathcal{U}\times\mathcal{A}$.
Algorithm III-B3 shows the procedure to compute the Stackelberg equilibrium for this class, comprising a game problem for the leaders and a centralized control problem for the followers.
@@ III-B4 Class IV: Stackelberg game with cooperative leaders and non-cooperative followers
The last possible combination of strategic behaviors consists of allowing the leaders to cooperate, meaning that this layer game problem is solved by means of solving an optimization problem. Then, given that the followers $\mathcal{M}$ do not cooperate, this layer is solved by means of a best-response dynamics in the context of dynamic games. The following shows the required steps for the computation of the Stackelberg equilibrium:
  • For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the Nash equilibrium $\mathbf{u}^{*}\in\mathcal{U}$, which is denoted by $\mathbf{u}^{*}(\mathbf{a})$, meaning that the equilibrium strategic profile depends on the strategic selection of the leader. The computed Nash equilibrium is a non-cooperative best response against the leaders’ strategies, i.e., $\mathbf{u}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{NC}}_{\mathcal{M}}(\mathbf{a})$.
  • Using the equilibrium strategic profile $\mathbf{u}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader cooperative problem can be stated only in terms of $\mathbf{a}$, and the joint optimization for the leaders can be computed, i.e., $\tilde{\mathbf{a}}^{*}\in\mathcal{A}$ can be found.
  • Then, using the cooperative solution for the leaders $\tilde{\mathbf{a}}^{*}$, the Nash equilibrium for the follower game problem becomes $\mathbf{u}^{*}:=\mathbf{u}^{*}(\tilde{\mathbf{a}}^{*})$.
  • Finally, one obtains the Class IV Stackelberg equilibrium as $(\mathbf{u}^{*},\tilde{\mathbf{a}}^{*})\in\mathcal{U}\times\mathcal{A}$.
Algorithm III-B4 shows the detailed procedure to compute the Stackelberg equilibrium for this class, comprising a dynamic game problem for the followers and an optimization problem for the leaders.
@@
After having introduced the theoretical foundation of the bi-level game problems for leaders and followers interacting in the framework of Stackelberg games, we present, in the following section, a networked system to illustrate an engineering application for the co-design problem.
@@ IV Networked System Application
The Barcelona Drinking Water Network (DWN) in [18, Chapter 3, Fig. 3.4], managed by the company Aguas de Barcelona (AGBAR), supplies drinking water to Barcelona city and its metropolitan area, utilizing water from the Ter and Llobregat rivers—regulated upstream by dams with a combined reservoir capacity of 600 hm³, alongside groundwater from the Besós River aquifer and supplementary wells. The network integrates four drinking water treatment plants: the Abrera and Sant Joan Despì facilities (Llobregat River), the Cardedeu plant (Ter River), and the Besós plant (groundwater), with additional pumping infrastructure extracting from wells to achieve a total flow of approximately 7 m³/s. This is a well-known benchmark that has been used to illustrate networked control applications, game-theory-based control analysis, and optimization-based controllers. Here, we use this case study to evaluate and illustrate the Stackelberg game classes introduced in this paper.
@@ IV-A Leaders and Followers in the Barcelona DWN
In terms of the co-design problem of the Barcelona DWN, we consider $L=4$ leaders within the system in charge of the design of four reservoirs’ dimensions, i.e., =$\mathcal{L}=\{1,\dots,4\}$ corresponding to the states $x_{1},x_{4},x_{10},x_{14}$, respectively (see Table II). Each leader in $\mathcal{L}$ has a set of possible actions in the strategic game
@@ 4 Red 2 $x_{1}$
On the other hand, let us consider $M=4$ followers, i.e., =$\mathcal{M}=\{1,\dots,4\}$, which are in charge of four different sub-systems of the DWN. As shown in Fig. 2, four subsystems highlighted by different colors are considered. This partitioning has been adopted from the research reported in [19]. Thus, each follower decides the control inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{4}$ corresponding to the sub-systems with colors green, blue, magenta, and red, respectively.
Note that the strategic selection of the followers $\mathcal{M}$ is influenced by the strategic selection of the leaders $\mathcal{L}$. When the leaders decide on a design for the reservoirs at each one of the sub-systems, the followers modify accordingly their optimal strategies. In the following, we introduce the game theoretical problems for both the leaders and followers, considering either cooperative or non-cooperative behavior. Let us start by introducing the normal form game problems for the leaders.
@@ IV-B Non-Cooperative and Cooperative Leaders Game
where $Q_{j}\succeq 0$ and $S_{ij}\succeq 0$ are weighting matrices, $v_{j}$ is a given vector. For leader costs, the parameters are chosen as $Q_{j}=0$, $v_{j}=0.01$, and $\alpha_{i,k}$ is the vector of time-varying electricity prices per input unit for follower $i$ at time $k$. The function $g(\mathbf{a})$ is used to penalize the effort that the leader applies in the design of the reservoirs. This cost can be associated with economic costs for implementing the design. In this regard, the leader is interested in minimizing the magnitude of its strategic selection. On the other hand, the cost function $h(\mathbf{u}$ depends on the followers’ strategic selection. This means that, when the leader makes decisions over the modeling, it also takes into consideration how the followers will perform their control actions. Note that this is a game problem as the cost functional of the $j-$th decision-maker (leader) is affected by the decisions made by $\mathcal{L}\setminus\{j\}$ through the followers’ actions $\mathbf{u}_{i}$, for all $i\in\mathcal{M}$. To emphasize this coupling, notice that the strategic design of a single leader, e.g., the $j-$th leader, affects the evolution of the control actions for all the followers $\mathcal{M}$ as they are dynamically coupled through (1). Therefore, as all the leaders $\mathcal{L}$ take into consideration all the followers’ actions in their cost functionals, the decisions of the $j-$th leader affects the cost of all the other leaders $\mathcal{L}\setminus\{j\}$ through $\mathbf{u}_{i}$ for all $i\in\mathcal{M}$.
@@ IV-C Non-Cooperative and Cooperative Followers Game
where $\Delta{u}_{i,k}={u}_{i,k}-{u}_{i,k-1}$ and $R_{i}\succ 0$, and $T>0$ is a planning horiizon. The feasible sets for both the control strategies and system states are as follows:
@@ V Results and Discussions
All simulations were conducted over a 72-hour period (3 days) with a sampling time interval of 1 hour. Water demands and electricity prices exhibit distinct daily patterns. The selected optimal solutions of the followers are shown in Fig. 3. In Figs. 3(a) and 3(b), the optimal follower solutions vary significantly, primarily attributable to differences in tank designs across four distinct classes and the implementation of two different control strategies (cooperative versus non-cooperative). Specifically, Fig. 3(b) highlights the scenarios where the inputs occasionally reach the maximum values because identical input constraints are applied. In Fig. 3(c), due to the same leader solutions with four classes, the same optimal follower solutions are observed. Fig. 3(d) further shows the impact of two control strategies: the follower solutions with Class I and Class IV, both governed by a non-cooperative control framework, exhibit striking similarities, as do those for Classes II and III, which adopt a cooperative control approach. Moreover, it can also be observed that the follower solutions, i.e., the flows through actuators (valves and pumps), have a potential daily pattern due to the water demand satisfaction.
@@
The optimal costs for both leaders and followers, and for all the possible classes (from Class I to Class IV), are presented in Table III and Table IV. By using such optimal values, one can measure or evaluate the difference between cooperating and non-cooperating, or from the control perspective, one can evaluate the cost difference of a centralized controller in front of a decentralized controller. We perform this assessment by means of the price-of-anarchy introduced in Section II. The results of the price-of-anarchy corresponding to all the interactive combinations are presented in Table V. It is interesting to observe that the price-of-anarchy is quite close to one in all the cases, indicating that the obtained Nash equilibria are optimal. In other words, the decentralized control approach design is optimal.

### v2

@@ III-B2 Class II: Stackelberg game with cooperative leaders and cooperative followers
Unlike Class I, where both leaders and followers act independently in a non-cooperative manner, Class II assumes full cooperation within each layer. Here, leaders jointly coordinate their strategic decisions, and followers collaboratively optimize their control actions. This cooperative interaction transforms the hierarchical game into a sequence of centralized optimization problems rather than nested non-cooperative games.
The procedure to compute the Stackelberg equilibrium in this cooperative setting is as follows:
  • Joint Follower Optimization: For each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, let us compute the optimal solution for the cooperative problem given by $\tilde{\mathbf{u}}^{*}\in\mathcal{U}$, which is denoted by $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ as it depends on the strategic selection of the leader. The computed optimal is a cooperative best response against the leaders’ strategies, i.e., $\tilde{\mathbf{u}}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{C}}_{\mathcal{M}}(\mathbf{a})$.
  • Joint Leader Optimization: Using the optimal strategic profile $\tilde{\mathbf{u}}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader cooperative game problem can be stated only in terms of $\mathbf{a}$, and the optimal solution for the leaders $\tilde{\mathbf{a}}^{*}\in\mathcal{A}$ can be found.
  • Stackelberg Equilibrium: Class II Stackelberg equilibrium can be found as $(\tilde{\mathbf{u}}^{*},\tilde{\mathbf{a}}^{*})\in\mathcal{U}\times\mathcal{A}$.
Class II replaces iterative best-response dynamics typical of non-cooperative games with a two-layer hierarchical optimization of coupled cost functions. This fundamental structural difference enables more tractable computations under cooperation assumptions, as outlined in Algorithm III-B2.
@@ III-B3 Class III: Stackelberg game with non-cooperative leaders and cooperative followers
Class III represents a mixed strategic setting in which the leaders $\mathcal{L}$ and followers $\mathcal{M}$ adopt fundamentally different behaviors. Specifically, the leaders behave independently, each optimizing their own individual cost function without cooperation within the leader layer. This structure leads to a two-layer approach: a non-cooperative game among leaders solved via best-response dynamics, and a cooperative control problem solved collectively by the followers. The interplay between these layers defines the Stackelberg equilibrium for this class.
The computation steps of the Stackelberg equilibrium are analogous to the ones in Class I. First, for each feasible leader strategic selection $\mathbf{a}\in\mathcal{A}$, the followers solve a centralized cooperative optimal control problem to obtain their best response $\tilde{\mathbf{u}}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{C}}_{\mathcal{M}}(\mathbf{a})$. Then, using this best response, the leaders engage in a non-cooperative game among themselves, where each optimizes its individual objective, leading to a Nash equilibrium $\mathbf{a}^{*}\in\mathcal{A}$. The followers then update their strategy accordingly to $\tilde{\mathbf{u}}^{*}:=\tilde{\mathbf{u}}^{*}(\mathbf{a}^{*})$. Together, the pair $(\tilde{\mathbf{u}}^{*},\mathbf{a}^{*})\in\mathcal{U}\times\mathcal{A}$ constitutes the Stackelberg equilibrium for this class, capturing the strategic interaction between non-cooperative leaders and cooperative followers.
Class III captures realistic scenarios where leaders may compete or have conflicting interests, while followers coordinate actions for mutual benefit. Algorithm III-B3 provides a detailed procedural framework for computing this equilibrium.
@@ III-B4 Class IV: Stackelberg game with cooperative leaders and non-cooperative followers
Class IV corresponds to the case where the leaders engage in cooperative behavior while the followers act in a non-cooperative, competitive manner. In this configuration, the leaders jointly solve an optimization problem to determine their collective optimal strategy, whereas the followers engage in a dynamic game, reaching equilibrium through best-response dynamics.
The computation of the Stackelberg equilibrium for this setting proceeds as follows. First, for every feasible leader strategic profile $\mathbf{a}\in\mathcal{A}$, the followers solve a non-cooperative dynamic game to compute the corresponding Nash equilibrium $\mathbf{u}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{NC}}_{\mathcal{M}}(\mathbf{a})$, which represents the followers’ best responses to the leaders’ decisions. Second, using the equilibrium strategic profile $\mathbf{u}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader cooperative problem can be stated only in terms of $\mathbf{a}$, and the joint optimization for the leaders can be computed, i.e., $\tilde{\mathbf{a}}^{*}\in\mathcal{A}$ can be found. Then, using the cooperative solution for the leaders $\tilde{\mathbf{a}}^{*}$, the Nash equilibrium for the follower game problem becomes $\mathbf{u}^{*}:=\mathbf{u}^{*}(\tilde{\mathbf{a}}^{*})$. The resulting pair $(\mathbf{u}^{*},\tilde{\mathbf{a}}^{*})$ constitutes the Class IV Stackelberg equilibrium, integrating the leaders’ cooperative optimization with the followers’ strategic competition. The complete procedure for this computation is detailed in Algorithm III-B4.
@@
III-C Computations for Four Classes
While the four classes of Stackelberg games provide a structured and systematic framework for modeling heterogeneous strategic interactions, the computation of their corresponding equilibria can be computationally demanding, especially for large-scale systems with multiple decision-makers. The primary source of this computational burden stems from the necessity to solve either optimization problems or game-theoretic equilibrium problems.
For Classes I and IV, the followers’ layer involves the computation of Nash equilibria in a dynamic game setting. This typically requires the implementation of iterative best-response algorithms or fixed-point methods, where each follower repeatedly adjusts their strategy based on the actions of others. The convergence of these iterative methods may require a substantial number of iterations, particularly when the interaction among players is highly nonlinear or when their strategy spaces are large. Moreover, because this Nash equilibrium computation must be repeated for every feasible strategic profile of the leaders, the computational cost scales quickly with the number of players and the size of their strategy sets.
For Classes II and III, while the followers act cooperatively and thus their problem reduces to solving a centralized optimal control problem, the leaders’ layer (Class III) or both layers (Class II) may still involve solving large-scale optimization problems over potentially high-dimensional spaces. Even though cooperative optimization problems tend to be more tractable than general Nash equilibrium computations, they can still be computationally heavy if the problem involves nonlinear dynamics, long planning horizons, or complex constraints.
In this paper, we consider finite feasible options for the leaders’ strategic selections, which simplifies part of the computational process by enabling exhaustive search or combinatorial optimization techniques at the leader level. Additionally, the followers’ best-response dynamics can be computed for each leader selection.
III-D Convergence Discussions of Algorithms
The convergence of both best-response-based algorithms and optimization-based methods has been extensively studied in the literature, particularly for problems involving continuous, convex, and coercive cost functionals with respect to the decision variables (strategies), as we have stated in the respective problem statements, e.g., linear-quadratic difference games [26, 27, 28, 29].
However, in the context of the Stackelberg game structures considered here, additional challenges arise. Specifically, the function $\mathbf{u}(\mathbf{a})$, representing the optimal response of the followers to a given leader strategy $\mathbf{a}$, introduces significant complexity. This makes the leader-level objective functions, $J_{j}(a_{j},a_{-j},\mathbf{u}(\mathbf{a}))$, $\forall j\in\mathcal{L}$, generally non-convex and difficult to handle analytically. Unlike standard Stackelberg game formulations where leaders and followers are coupled through shared dynamics or additive costs (e.g., [6, 24, 25]), here the leaders’ strategies $\mathbf{a}$ affect the constraints of the follower optimization problem $\mathcal{P}_{\mathrm{follower}}$.
To address this challenge, the proposed algorithms compute a numerical approximation of the optimal follower response $\mathbf{u}(\mathbf{a})$ for each possible leader strategy $\mathbf{a}$. By working directly with these numerical solutions, the leader-level cost function $J_{j}(a_{j},a_{-j},\mathbf{u}(\mathbf{a}))$ can be structured to retain desirable properties, such as convexity with respect to $a_{j}$. This numerical decoupling simplifies the problem structure without requiring explicit expressions for $\mathbf{u}(\mathbf{a})$.
Provided that the follower cost functionals $V_{i}(x_{0},\mathbf{u})$ are continuous, convex, and coercive with respect to each follower’s control $\mathbf{u}i$, and that the feasible set of $\mathcal{P}{\mathrm{follower}}$ is non-empty for each $\mathbf{a}$, existence of a solution to the follower subproblem is guaranteed. Similarly, given a numerically computed $\mathbf{u}^{*}$, the leader optimization problem also admits a solution under the assumed continuity, convexity, and coerciveness of $J_{j}$ with respect to each $a_{j}$.
@@ IV Networked System Application
In this section, we apply the proposed four classes in a real-world drinking water network. This case study demonstrates that the proposed framework can produce co-design solution for large-scale networked systems. Additionally, this framework can also be extended to other networked systems, such as power grids, transportation networks, and supply chains.
The Barcelona Drinking Water Network (DWN) in [30, Chapter 3, Fig. 3.4], managed by the company Aguas de Barcelona (AGBAR), supplies drinking water to Barcelona city and its metropolitan area, utilizing water from the Ter and Llobregat rivers—regulated upstream by dams with a combined reservoir capacity of 600 hm³, alongside groundwater from the Besós River aquifer and supplementary wells. The network integrates four drinking water treatment plants: the Abrera and Sant Joan Despì facilities (Llobregat River), the Cardedeu plant (Ter River), and the Besós plant (groundwater), with additional pumping infrastructure extracting from wells to achieve a total flow of approximately 7 m³/s. This is a well-known benchmark that has been used to illustrate networked control applications, game-theory-based control analysis, and optimization-based controllers. Here, we use this case study to evaluate and illustrate the Stackelberg game classes introduced in this paper.
@@ IV-A Leaders and Followers in the Barcelona DWN
In terms of the co-design problem of the Barcelona DWN, we consider $L=4$ leaders within the system in charge of the design of four reservoirs’ dimensions, i.e., =$\mathcal{L}=\{1,\dots,4\}$ corresponding to the states $x_{4},x_{10},x_{14},x_{1}$, respectively (see Table II). Each leader in $\mathcal{L}$ has a set of possible actions in the strategic game
@@ 4 Red 2 $x_{1}$
On the other hand, let us consider $M=4$ followers, i.e., =$\mathcal{M}=\{1,\dots,4\}$, which are in charge of four different sub-systems of the DWN. As shown in Fig. 2, four subsystems highlighted by different colors are considered. This partitioning has been adopted from the research reported in [30, Chapter 8, Fig. 8.3]. Thus, each follower decides the control inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{4}$ corresponding to the sub-systems with colors green, blue, magenta, and red, respectively.
Note that the strategic selection of the followers $\mathcal{M}$ is influenced by the strategic selection of the leaders $\mathcal{L}$. When the leaders decide on a design for the reservoirs at each one of the sub-systems, the followers modify accordingly their optimal strategies. In the following, we introduce the game-theoretic problems for both the leaders and followers, considering either cooperative or non-cooperative behavior. Let us start by introducing the normal form game problems for the leaders.
@@ IV-B Non-Cooperative and Cooperative Leaders Game
where $Q_{j}\succeq 0$ is weighting matrices, $v_{j}$ is a given vector. For leader costs, the parameters are chosen as $Q_{j}=\mathbf{0}$ (inspired by an economic cost shape), $v_{j}=0.01$, and $\alpha_{i,k}$ is the vector of time-varying electricity prices per input unit for follower $i$ at time $k$. The function $g(\mathbf{a})$ is used to penalize the effort that the leader applies in the design of the reservoirs. This cost can be associated with economic costs for implementing the design. In this regard, the leader is interested in minimizing the magnitude of its strategic selection. On the other hand, the cost function $h(\mathbf{u}$ depends on the followers’ strategic selection. This means that, when the leader makes decisions over the modeling, it also takes into consideration how the followers will perform their control actions. Note that this is a game problem as the cost functional of the $j-$th decision-maker (leader) is affected by the decisions made by $\mathcal{L}\setminus\{j\}$ through the followers’ actions $\mathbf{u}_{i}$, for all $i\in\mathcal{M}$. To emphasize this coupling, notice that the strategic design of a single leader, e.g., the $j-$th leader, affects the evolution of the control actions for all the followers $\mathcal{M}$ as they are dynamically coupled through (1). Therefore, as all the leaders $\mathcal{L}$ take into consideration all the followers’ actions in their cost functionals, the decisions of the $j-$th leader affects the cost of all the other leaders $\mathcal{L}\setminus\{j\}$ through $\mathbf{u}_{i}$ for all $i\in\mathcal{M}$.
@@ IV-C Non-Cooperative and Cooperative Followers Game
where $\Delta{u}_{i,k}={u}_{i,k}-{u}_{i,k-1}$ and $R_{i}\succ 0$, and $T>0$ is a planning horizon. The feasible sets for both the control strategies and system states are as follows:
@@ V Results and Discussions
All simulations for followers were conducted over a 72-hour period (3 days) with a sampling time interval of 1 hour. Water demands and electricity prices exhibit distinct daily patterns. The selected optimal solutions of the followers are shown in Fig. 3. In Figs. 3(a) and 3(b), the optimal follower solutions vary significantly, primarily attributable to differences in tank designs across four distinct classes and the implementation of two different control strategies. Specifically, Fig. 3(b) highlights the scenarios where the inputs occasionally reach the maximum values because identical input constraints are applied. In Fig. 3(c), due to the same leader solutions with four classes, the same optimal follower solutions are observed. Fig. 3(d) further shows the impact of two control strategies: the follower solutions with Class I and Class IV, both governed by a non-cooperative control framework, exhibit striking similarities, as do those for Classes II and III, which adopt a cooperative control approach. Moreover, it can also be observed that the follower solutions, i.e., the flows through actuators (valves and pumps), have a potential daily pattern due to the water demand satisfaction.
@@
The optimal costs for both leaders and followers, and for all the possible classes (from Class I to Class IV), are presented in Table III and Table IV. By using such optimal values, one can measure or evaluate the difference between cooperating and non-cooperating, or from the control perspective, one can evaluate the cost difference of a centralized controller in front of a decentralized controller. We perform this assessment by means of the price-of-anarchy introduced in Section II. The results of the price-of-anarchy corresponding to all the interactive combinations are presented in Table V. It is interesting to observe that the price-of-anarchy is quite close to one in all the cases, indicating that the obtained Nash equilibria are optimal.